OpenStudy (anonymous):

You roll 2 dice. What is the probability that the sum of the dice is greater than 7 and 1 die shows a 2? A 6 X 6 table of dice outcomes will help you to answer this question. Write fractions in lowest terms.

OpenStudy (anonymous):

@terenzreignz

OpenStudy (anonymous):

count!!

OpenStudy (anonymous):

(2,6) works so does (6,2) i think that is all

OpenStudy (anonymous):

Total number of possible outcomes = 6^2 = 36 Sum of the digits---No. of outcomes 02 ------------------------- 01 03 ------------------------- 02 04 ------------------------- 03 05 ------------------------- 04 06 ------------------------- 05 07 ------------------------- 06 08 ------------------------- 05 09 ------------------------- 04 10 ------------------------- 03 11 ------------------------- 02 12 ------------------------- 01 Total --------------------- 36 P(sum of the digits is odd) = P(A) = (2+4+6+4+2) / 36 = 18/36 P(sum of the digits > 6) = P(B) = (6+5+4+3+2+1) / 36 = 21/36 P(A and B) = (6+4+2) / 36 = 12/36 Required probability = P(A or B) = P(A)+P(B)-P(A and B) =18/36 + 21/36 - 12/36 = 27/36 = 3/4

OpenStudy (anonymous):

lol some other answer to some other question

OpenStudy (anonymous):

that doesn't help:(

OpenStudy (anonymous):

2,6 & 6,2 aren't the only ones

OpenStudy (anonymous):

just count is all how many fit the description

OpenStudy (anonymous):

First notice that the sample space contains 6 x 6 = 36 possibilities. Now notice that our event space only contains those possibilities that show a 2 and have a sum greater than 7. The only way to have a sum greater than 7 is when 2 + x > 7, where x is some other number we rolled. Solve the inequality and we get:\[\bf 2+x>7 \implies x > 5\]The only other number we can roll which is greater than 5 is 6. Hence the two ways we can roll 2 dice and get a 2 and a sum greater than 7 is (2,6) or (6,2). To computer the probability, we know there is 36 possible ways and only 2 ways that we want. \[\bf P(\sum > 7 \,, shows \ 2)= \frac{ n(A) }{ n(S) }=\frac{ 2 }{ 36 }=??\]

OpenStudy (anonymous):

ok find another and show that i am not right

OpenStudy (anonymous):

@music101

OpenStudy (anonymous):

oh it has to be GREATER than huh? not equal to 7???

OpenStudy (anonymous):

right!!

OpenStudy (anonymous):

so itd be: 4,4 5,4 4,5 6,5 5,6

OpenStudy (anonymous):

oh hell no

OpenStudy (anonymous):

don't forget is says AND one die shows "2"

OpenStudy (anonymous):

nnooooooooooo @music101 look at what I just said. There MUST BE A 2 and the sum MUST BE GREATER THAN 7

OpenStudy (anonymous):

oh oops

OpenStudy (anonymous):

@satellite73 I must say, you've readily improved at probability questions.

OpenStudy (anonymous):

so it is just 2,6 & 6,2 then

OpenStudy (anonymous):

YES!!!!

OpenStudy (anonymous):

So 2 ways to get a 2 and a sum greater than 7 out of a total of 36 possible ways. So the probability is??? @music101

OpenStudy (anonymous):

1/18?

OpenStudy (anonymous):

YESESESE@!#!@# U DID IT!

OpenStudy (anonymous):

thank u.. sorry probability is just a little confusing.. ugh, there were worse problems

OpenStudy (anonymous):

I love them. If you got more questions, just post them and I'll help you out. @musicalcat98

OpenStudy (anonymous):

ok ill post them now

OpenStudy (anonymous):

who's musicalcat??

OpenStudy (ybarrap):

I think that probability is just nature's sinister way of showing mathematician's that they can't even count. Counting is the hardest part of probability, I think.